# Identifying faulty edges in resistive electrical networks

### v1

@@ Abstract
Given a resistive electrical network, we would like to determine whether all the resistances (edges) in the network are working, and if not, identify which edge (or edges) are faulty. To make this determination, we are allowed to measure the effective resistance between certain pairs of nodes (which can be done by measuring the amount of current when one unit of voltage difference is applied at the chosen pair of nodes). The goal is to determine which edge, if any, is not working in the network using the smallest number of measurements. We prove rigorous upper and lower bounds on this optimal number of measurements for different classes of graphs. These bounds are tight for several of these classes showing that our measurement strategies are optimal.
@@
1 Background and Motivation
Diagnosing faults in electrical circuits and wiring is an important problem with applications in power systems, very large scale integration (VLSI), chip design, circuit board systems, to name a few. The literature is vast; the reader is referred to the survey [furse2020fault], a sample of more recent work [dwivedi2010fault, wang2019electromagnetic] and the references therein. There are a variety of approaches that are adopted in practice and new methods are regularly being invented even to the present. In this paper, we consider a particular formalization of this problem which, to the best of our knowledge, has not been studied rigorously in prior work. Nevertheless, we believe the model we study is natural, of practical relevance, and worth studying from a mathematical perspective. It is related to, and inspired by, work in detecting faults and structures in continuous conducting media; see the textbooks [holder2004electrical, bui2007fracture] and [adler2015electrical] for a more recent survey. Our problem can be seen as a discrete analogue of this problem from the literature on continuous media and related inverse problems. The resulting optimization problem is rich in its scope, spanning topics from combinatorial and spectral graph theory, integer and combinatorial optimization, information theory, and electrical network theory. Let us formally introduce the problem.
@@ 1 Background and Motivation
Let $\mathcal{A}\in\mathbb{R^{|T|x|E|}}$ be the matrix where each row corresponds to a measurement $t\in T$ and each column corresponds to an edge $e\in E$ that has been altered. Each entry in $\mathcal{A}$ contains the effective resistance value for a given pair of source and sink terminals and a given altered edge. Consider the submatrix $\mathcal{A}_{S}\in\mathbb{R^{|S|x|E|}}$ of values corresponding to measurements $S\subseteq T$ and all edges $E$. If all columns in $\mathcal{A}_{S}$ are different from one another, then the provided measurements $S$ satisfy the version of our problem given in Definition 1.2. On the other hand, if there are at least two columns that are the same, then the set $S$ of measurements will not distinguish between these two edges. Hence, an equivalent formulation of the problem is to find the smallest set $S\subseteq T$ of measurements such that $\mathcal{A}_{S}$ has distinct columns.
@@ 1.1 Our contributions
We provide tight upper and lower bounds on the smallest number of measurements for the fault detection problem in various families of electrical networks. In particular, we are able to fully resolve the problem for complete graphs and complete $k$-partite graphs for $k\geq 2$. The formal statements are in Section 2 and their proofs are provided in Section LABEL:sec:proofs. The proof techniques draw upon combinatorial and spectral graph theory, electrical network theory, and ideas from information theory. We note briefly that each effective resistance measurement provides a real number; thus, a single measurement potentially provides infinitely many bits of information. For this reason, classical information theoretic lower bound arguments based on bits of information do not immediately apply. While our final results are mathematically precise and rigorous, our search for the best measurement strategy in each of these cases involved modeling the problem as a set covering/integer programming problem and using computational experiments to guide us towards the optimal strategy for measurements. Thus, the problem involves a nice interplay between rigorous mathematical analysis and computer-aided search using discrete optimization techniques. More broadly, we believe that the fault detection problem we set up in Definition 1.2 is a new, challenging combinatorial optimization problem that is rich in its mathematical structure and is motivated by a fundamental problem in fault diagnosis in electrical networks, which has diverse applications. We close the paper in Section LABEL:sec:future_avenues with some future directions, including some concrete open questions and conjectures, that we hope can spawn new and interesting research avenues for the combinatorial and discrete optimization community.
@@ 1.2 Related literature
In [shi1999diagnosis, shi2001structural], the authors consider the problem of detecting faults in wiring networks, building on prior work which can be traced back to a seminal paper by Kautz [kautz1974testing]. Here, the vertices of a graph model different logical units (often called nets) in a circuit board and each edge denotes a possible site of a “fault”. One is allowed to probe the logical units and use the responses to detect these faults, if any. The mathematical formalization is the following: given a (simple) graph $G=(V,E)$, one can query a subset $U\subseteq V$ and the response is a subset $Q(U)$ which is the set of all vertices that can be reached by a path from a vertex in $U$. The goal is to find the smallest set of queries to identify all the connected components of $G$ (the vertices in the connected components of size two or more correspond to the logical units that are connected by faulty edges). A group testing [aldridge2019group] approach to this problem was considered in [chen1989detecting] and is discussed in the papers [shi1999diagnosis, shi2001structural] cited above, along with other relevant work on this problem.
@@ 2 Statement of results
Complete graphs.
The following theorem gives the number of measurements needed for complete graphs (i.e., graphs that have an edge between every pair of vertices).
@@
Complete $k$-partite graphs.
@@
When $|p_{\beta}|<|p_{\gamma}|$, the smallest number of measurements needed to solve the faulty edge detection problem is $\left\lfloor\frac{2}{3}|p_{\gamma}|+\frac{1}{3}|p_{\beta}|\right\rfloor-1$ when $|p_{\gamma}|-|p_{\beta}|\equiv 0\,(\operatorname{mod}3)$, and equal to $\left\lfloor\frac{2}{3}|p_{\gamma}|+\frac{1}{3}|p_{\beta}|\right\rfloor$ when $|p_{\gamma}|-|p_{\beta}|\equiv 1\textrm{ or }2\,(\operatorname{mod}3)$.
When $|p_{\beta}|=|p_{\gamma}|$, the smallest number of measurements needed to solve the faulty edge detection problem is $\left\lfloor\frac{2n}{3}\right\rfloor-1$ when $|p_{\beta}|\equiv 0\textrm{ or }1\,(\operatorname{mod}3)$, and equal to $\left\lfloor\frac{2n}{3}\right\rfloor$ when $|p_{\beta}|\equiv 2\,(\operatorname{mod}3)$.

### v2

@@ Abstract
Given a resistive electrical network, we would like to determine whether all the resistances (edges) in the network are working, and if not, identify which edge (or edges) is faulty. To make this determination, we are allowed to measure the effective resistance between certain pairs of nodes (which can be done by measuring the amount of current when one unit of voltage difference is applied at the chosen pair of nodes). The goal is to determine which edge, if any, is not working in the network using the smallest number of measurements. We prove rigorous upper and lower bounds on this optimal number of measurements for different classes of graphs. These bounds are tight for several of these classes showing that our measurement strategies are optimal.
@@
1 Introduction
Diagnosing faults in electrical circuits and wiring is an important problem with applications in power systems, very large scale integration (VLSI), chip design, circuit board systems, to name a few. The literature is vast; the reader is referred to the survey [furse2020fault], a sample of more recent work [dwivedi2010fault, wang2019electromagnetic] and the references therein. There are a variety of approaches that are adopted in practice and new methods are regularly being invented even to the present. In this paper, we consider a particular formalization of this problem which, to the best of our knowledge, has not been studied rigorously in prior work. Nevertheless, we believe the model we study is natural, of practical relevance, and worth studying from a mathematical perspective. It is related to and inspired by work in detecting faults and structures in continuous conducting media; see the textbooks [holder2004electrical, bui2007fracture] and [adler2015electrical] for a more recent survey. Our problem can be seen as a discrete analogue of this problem from the literature on continuous media and related inverse problems. The resulting optimization problem is rich in its scope, spanning topics from combinatorial and spectral graph theory, integer and combinatorial optimization, information theory, and electrical network theory. Let us formally introduce the problem.
@@ 1 Introduction
Let $\mathcal{A}\in\mathbb{R^{|T|\times|E|}}$ be the matrix where each row corresponds to a measurement $t\in T$ and each column corresponds to an edge $e\in E$ that has been altered. Each entry in $\mathcal{A}$ contains the effective resistance value for a given pair of source and sink terminals and a given altered edge. Consider the submatrix $\mathcal{A}_{S}\in\mathbb{R^{|S|\times|E|}}$ of values corresponding to measurements $S\subseteq T$ and all edges $E$. If all columns in $\mathcal{A}_{S}$ are different from one another, then the provided measurements $S$ satisfy the version of our problem given in Definition 1.2. On the other hand, if there are at least two columns that are the same, then the set $S$ of measurements will not distinguish between these two edges. Hence, an equivalent formulation of the problem is to find the smallest set $S\subseteq T$ of measurements such that $\mathcal{A}_{S}$ has distinct columns.
@@
We also note that in Definition 1.2, the measurements $S\subseteq T$ are made non-adaptively, i.e., the result of a measurement is not used to decide what the next measurement is. One must decide on the set $S$ of measurements up front and then the observed measurement values must be used to identify the faulty edge. Our results below apply to this non-adaptive setting. The analysis of adaptive measurement strategies is left for future work.
@@ 1.1 Our contributions
We provide tight upper and lower bounds on the smallest number of measurements for the fault detection problem in various families of electrical networks. In particular, we are able to fully resolve the problem for complete graphs and complete $k$-partite graphs for $k\geq 2$. The formal statements are in Section 2 and their proofs are provided in Section LABEL:sec:proofs. The proof techniques draw upon combinatorial and spectral graph theory, electrical network theory, and ideas from information theory. We note briefly that each effective resistance measurement provides a real number; thus, a single measurement potentially provides infinitely many bits of information. For this reason, classical information theoretic lower bound arguments based on bits of information do not immediately apply. While our final results are mathematically precise and rigorous, our search for the best measurement strategy in each of these cases involved modeling the problem as a set covering/integer programming problem and using computational experiments to guide us towards the optimal strategy for measurements. Thus, the problem involves a nice interplay between rigorous mathematical analysis and computer-aided search using discrete optimization techniques. More broadly, we believe that the fault detection problem we set up in Definition 1.2 is a new, challenging combinatorial optimization problem that is rich in its mathematical structure and is motivated by a fundamental problem in fault diagnosis in electrical networks, which has diverse applications. We close the paper in Section LABEL:sec:future_avenues with some future directions, including some concrete open questions, that we hope can spawn new and interesting research avenues for the discrete mathematics and optimization community.
@@ 1.2 Related literature
The closest line of work is the investigation in [kahng1998test], which is motivated by a fault detection problem in VLSI systems. In this paper, the authors consider the problem of testing whether all edges in a tree network are working properly; thus, the problem is called the tree testing problem. In the language of Definition 1.2, the graph $G$ is a tree, all edge weights are $1$, and the goal is to simply detect if any edge has been altered or not, without necessarily pinpointing which one if a fault is indeed detected.
@@
In [shi1999diagnosis, shi2001structural], the authors consider the problem of detecting faults in wiring networks. They build on prior work which can be traced back to a seminal paper by Kautz [kautz1974testing]. Here, the vertices of a graph model different logical units (often called nets) in a circuit board and each edge denotes a possible site of a “fault”. One is allowed to probe the logical units and use the responses to detect these faults, if any. The mathematical formalization is the following: given a (simple) graph $G=(V,E)$, one can query a subset $U\subseteq V$ and the response is a subset $Q(U)$ which is the set of all vertices that can be reached by a path from a vertex in $U$. The goal is to find the smallest set of queries to identify all the connected components of $G$ (the vertices in the connected components of size two or more correspond to the logical units that are connected by faulty edges). A group testing [aldridge2019group] approach to this problem was considered in [chen1989detecting] and is discussed in the papers [shi1999diagnosis, shi2001structural] cited above, along with other relevant work on this problem.
@@ 2 Statement of results
Complete graphs
The following theorem gives the number of measurements
needed for complete graphs (i.e., graphs that have an edge between every pair of vertices).
@@
Complete $k$-partite graphs
@@
When $|p_{\beta}|<|p_{\gamma}|$, the smallest number of measurements needed to solve the faulty edge detection problem is $\left\lfloor{\frac{2}{3}|p_{\gamma}|+\frac{1}{3}|p_{\beta}|}\right\rfloor-1$ when $|p_{\gamma}|-|p_{\beta}|\equiv 0\mod{3}$, and equal to $\left\lfloor\frac{2}{3}|p_{\gamma}|+\frac{1}{3}|p_{\beta}|\right\rfloor$ when $|p_{\gamma}|-|p_{\beta}|\equiv(1\textrm{ or }2)\mod{3}$.
When $|p_{\beta}|=|p_{\gamma}|$, the smallest number of measurements needed to solve the faulty edge detection problem is $\left\lfloor\frac{2n}{3}\right\rfloor-1$ when $|p_{\beta}|\equiv 0\textrm{ or }1\mod{3}$, and equal to $\left\lfloor\frac{2n}{3}\right\rfloor$ when $|p_{\beta}|\equiv 2\mod{3}$.